Please add the computation and explanation. Thank you!
A study on the voting patterns of members of the Philippine Congress on Cha Cha issue is shown below: PARTY AFFILIATION NATIONALISTA LAKAS/OTHERS LIBERAL 60 82 70 For 69 Against Undecided 95 60 21 20 23 Test the Null Hypothesis at 0.05 level of significance that voting patterns do not depend on party affiliation. Tabulated below are the percentages of votes predicted by a survey conducted by a private agency for twelve senatorial candidates x and the percentages of votes which these candidates actually received, y during the 2016 elections. Calculate the coefficient of correlation and interpret the results. II Survey x Election y 54 49 49 47 30 28 33 29 47 40 50 44 36 38 25 20 18 19 27 24 30 28 43 41 I A.) Survey on the "Leadership Behavior" of 15 managers was conducted. Twenty-five employee respondents were taken. The results are found below using an instrument with how frequently the manager engages in the behavior described by them 5 4 3 2 1 Always Often Occasionally Seldom Never (Occ) 3 (A) 4 (Oft) 13 (N) (S) 1. He makes his attitude clear to the group 2. He is easy to understand. 3. He schedules the work to be done 3 2 2 3 12 6 2 5 13 3 4. He maintains definite standards of performance. 5. He is willing to make changes 4 2 6 12 2 2 3 2 16 N N en 3. 10
Question: 1. How did the manager behave in each statement as perceived by the employee respondents? Determine also the standard deviation. 2. B.) The same instrument was administered by the 15 managers and resulted into the data below. 1 3 Always Often Occasionally Seldom Never (A) (N) (Oft) (Occ) 5 (S) 2 1 1. He makes his attitude clear to the group 2. He is easy to understand 3. He schedules the work to be done 4. He maintains definite 3 4 2 1 2 7 3 2 2 5 2 4 2 1 1 7 4 standards of performance. 5. He is willing to make changes 2 2 8 1 2 Questions 1.) How did the managers rate themselves in each of the statement? 2.) Compute also the standard deviation Consider the sales of 4 stores. At 0.05 level of significance, is there a significant difference among the sales of the 4 stores? Follow the steps in hypothesis testing IV Store A Store C Store B Store D 25 18 12 25 28 20 22 26 27 18 21 21 12 23 11 23 LC
Answer #1

I

H0: Null Hypothesis: Voting patterns do not depend on party affiliation.

HA: Alternative Hypothesis: Voting patterns depend on party affiliation

Observed Frequencies:

LIBERAL NATIONALIST A LAKAS/ OTHERS Total
For 82 70 60 212
Against 95 60 69 224
Undecided 23 20 21 64
Total 200 150 150 500

Assuming H0, Expected Frequencies are calculated as follows:

Expected Frequencies:

LIBERAL NATIONALIST A LAKAS/ OTHERS Total
For 200X212/500=84.80 63.60 63.60 212
Against 89.60 67.20 67.20 224
Undecided 25.60 19.20 19.60 64
Total 200 150 150 500

The Test Statistic \chi ^{2} is calculated as follows:

Observed (O) Expected (E) (O - E)2/E
82 84.80 0.09
70 63.60 0.20
95 89.60 0.33
60 67.20 0.77
69 67.20 0.05
23 25.60 0.26
20 19.20 0.03
21 19.20 0.17
Total = \chi ^{2} = 2.55

ndf = (r - 1) X (c - 1)

   = (3 - 1) X (3 - 1)

   = 4

By Technology, P - Value = 0.6354

Since P - Value is greater than \alpha = 0.05, the difference is not significant. Fail to reject null hypothesis.

Conclusion:
The data support the claim that voting patterns do not depend on party affiliation.

AS PER DIRECTIVE, FIRST FULL QUESTION IS ANSWERED.